OpenStudy (anonymous):

y = 3x + 3b for b How do I solve for the indicated variable?

jimthompson5910 (jim_thompson5910):

\[\large y = 3x + 3b\] \[\large y - 3x = 3b\] \[\large \frac{y - 3x}{3} = b\] \[\large b=\frac{y - 3x}{3}\]